# Moduli spaces of untwisted wild Riemann surfaces

### v1

@@ 2.2.3. Properties of $\mathscr{T}_{\Sigma,\sigma}^{\leq k}$
For every $x\in\Sigma\smallsetminus\sigma(B)$ there is an open neighbourhood $U$ of $x$ which does not intersect $\sigma(B)$, hence $\mathcal{I}_{\sigma}|U=\mathcal{O}_{\Sigma}|U$. It follows that $\mathscr{T}_{\Sigma,\sigma}^{\leq k}$ is a sheaf supported on $\sigma(B)$ and killed by $I_{\sigma}^{k}$, so it is the pushforward of a unique sheaf, denoted by $\mathscr{T}_{\sigma}^{\leq k}$, on the complex analytic space $(\sigma(B),\mathcal{O}_{\Sigma}/I_{\sigma}^{k})$. In particular, even though $\mathscr{T}_{\Sigma,\sigma}^{\leq k}$ was defined as a quotient of a sheaf of meromorphic functions on the whole $\Sigma$, it only carries information about germs of such functions along $\sigma(B)$. Precisely, if $V\subset U\subset\Sigma$ are two opens having the same intersection with $\sigma(B)$, then $\mathscr{T}_{\Sigma,\sigma}^{\leq k}(U)=\mathscr{T}_{\Sigma,\sigma}^{\leq k}(V)$.
We will denote by $B^{(k)}$ the analytic space $(\sigma(B),\mathcal{O}_{\Sigma}/I_{\sigma}^{k})$, and by $\pi^{(k)}\colon B^{(k)}\rightarrow B$ the composition of $\pi$ and the inclusion of $B^{(k)}$ in $\Sigma$.
@@
The sheaf $\pi_{*}\mathscr{T}_{\Sigma,\sigma}^{\leq k}$ coincides with the pushforward of $\mathscr{T}_{\sigma}^{\leq k}$ via the map $\pi^{(k)}\colon B^{(k)}\rightarrow B$. Choosing local coordinates around a point of $\sigma(B)$ as in section 2.1.2, we see that we can write locally $\mathscr{T}_{\Sigma,\sigma}^{\leq k}=z_{n+1}^{-k}\mathcal{O}_{\Sigma}/\mathcal{O}_{\Sigma}$, hence $\mathscr{T}_{\sigma}^{\leq k}$ is locally free of rank one. On the other hand, we claim that $\pi^{(k)}_{*}\mathcal{O}_{B^{(k)}}$ is a locally free $\mathcal{O}_{B}$-module of rank $k$. To prove this, we may, and will, replace $B$ by an open neighbourhood of any of its points, and work with the same local coordinates as before. The section $\sigma:B\rightarrow\Sigma$ factors through $B^{(k)}$, and it induces the natural projection $\mathcal{O}_{\Sigma}/z_{n+1}^{k}\mathcal{O}_{\Sigma}\rightarrow\mathcal{O}_{\Sigma}/z_{n+1}\mathcal{O}_{\Sigma}\simeq\mathcal{O}_{B}$. The map $\pi^{(k)}$ induces on global sections a section of this projection; hence, the claim follows from Lemma 2.2.5 below. ∎
@@
Let $A$ be a ring, $f\in A$ a nonzerodivisor, $k\geq 1$ an integer and $s\colon A/(f)\rightarrow A/(f^{k})$ a section of the projection $q\colon A/(f^{k})\rightarrow A/(f)$. The map $s$ endows $A/(f^{k})$ with the structure of a free $A/(f)$-module of rank $k$ with basis $1,f,\ldots,f^{k-1}$.
@@
By induction on $k$; for $k=1$ the claim is clear, so let us suppose that $k\geq 2$. Take $a\in A$ and let $\alpha_{0}=q(a)$. Then $q(a-s(\alpha_{0}))=q(a)-q\circ s\circ q(a)=0$, so we can write $a=s(\alpha_{0})+fa_{1}$ for some $a_{1}\in A/(f^{k})$. Assume that $\alpha_{0}^{\prime}\in A/(f),a_{1}^{\prime}\in A/(f^{k})$ also satisfy $a=s(\alpha_{0}^{\prime})+fa_{1}^{\prime}$. Then $s(\alpha_{0}-\alpha_{0}^{\prime})=f(a_{1}^{\prime}-a_{1})$; taking the image via $q$ we find that $\alpha_{0}=\alpha_{0}^{\prime}$, hence $f(a_{1}^{\prime}-a_{1})=0\in A/(f^{k})$. As $f$ is not a zero divisor, we deduce that $a_{1}^{\prime}-a_{1}\in(f^{k-1})$. Therefore, we have proved that $\alpha_{0}$ is unique, and $a_{1}$ is unique in $A/(f^{k-1})$, which implies the statement by induction. ∎
@@ 2.2.6. Functoriality
Let us set $\mathscr{T}_{|B}^{\leq k}=\pi_{*}\mathscr{T}_{\Sigma,\sigma}^{\leq k}$. Let $\varphi\colon B^{\prime}\rightarrow B$ be a holomorphic map, $\pi^{\prime}\colon\Sigma^{\prime}\rightarrow B^{\prime}$ the base change of $\pi$, and $\sigma^{\prime}\colon B^{\prime}\rightarrow\Sigma^{\prime}$ the pullback of $\sigma$. We have a commutative diagram
@@
and $\sigma^{\prime}(B^{\prime})\subset\Sigma^{\prime}$ is cut out by the ideal sheaf image of $\psi^{*}\mathcal{I}_{\sigma}$; in other words, we have $\sigma^{\prime}_{*}\mathcal{O}_{B^{\prime}}=\mathcal{O}_{\Sigma^{\prime}}/\psi^{*}\mathcal{I}_{\sigma}=\psi^{*}\sigma_{*}\mathcal{O}_{B}$.
@@
The following assertions hold true.
@@
    $\psi^{*}\mathcal{I}_{\sigma}=\mathcal{I}_{\sigma^{\prime}}$.
@@
    $\psi^{*}\mathcal{O}_{\Sigma}(\sigma(B))=\mathcal{O}_{\Sigma^{\prime}}(\sigma^{\prime}(B^{\prime}))$.
@@
    Pulling back the exact sequence $0\rightarrow\mathcal{I}_{\sigma}\rightarrow\mathcal{O}_{\Sigma}\rightarrow\sigma_{*}\mathcal{O}_{B}\rightarrow 0$ via $\psi$, and using that $\psi^{*}\sigma_{*}\mathcal{O}_{B}=\sigma^{\prime}_{*}\mathcal{O}_{B^{\prime}}$ we get an exact sequence
@@
    Pulling back the identity $\mathcal{I}_{\sigma}\otimes_{\mathcal{O}_{\Sigma}}\mathcal{O}_{\Sigma}(\sigma(B))\simeq\mathcal{O}_{\Sigma}$ and using (1) we deduce that $\psi^{*}(\mathcal{O}_{\Sigma}(\sigma(B)))$ is the inverse of $\mathcal{I}_{\sigma^{\prime}}$, hence it coincides with $\mathcal{O}_{\Sigma^{\prime}}(\sigma^{\prime}(B^{\prime}))$.
@@
    The $k$-fold tensor product of the isomorphism in (2) yields an isomorphism $\psi^{*}\mathcal{O}_{\Sigma}(k\sigma(B))=\mathcal{O}_{\Sigma^{\prime}}(k\sigma^{\prime}(B^{\prime}))$, which implies the claim by right-exactness of pullback. ∎
By adjunction we have a natural transformation $\mathrm{Id}\rightarrow\psi_{*}\psi^{*}$, inducing $\pi_{*}\rightarrow\pi_{*}\psi_{*}\psi^{*}$, so $\pi_{*}\rightarrow\varphi_{*}\pi^{\prime}_{*}\psi^{*}$ and finally $\varphi^{*}\pi_{*}\rightarrow\pi^{\prime}_{*}\psi^{*}$. Applying these functors to $\mathscr{T}_{\Sigma,\sigma}^{\leq k}$ and using the previous lemma we get a map $\varphi^{*}\mathscr{T}_{|B}^{\leq k}\rightarrow\mathscr{T}_{|B^{\prime}}^{\leq k}$.
@@
The above map $\varphi^{*}\mathscr{T}_{|B}^{\leq k}\rightarrow\mathscr{T}_{|B^{\prime}}^{\leq k}$ is an isomorphism.
@@
By [Sta18, Tag 01HQ] and Lemma 2.2.7 there is a unique map $\gamma\colon B^{\prime(k)}\rightarrow B^{(k)}$ making the top square of the diagram in section 2.2.6 Cartesian. We claim that the following diagram is Cartesian as well:
@@ 2.3. Irregular types
Let $\pi\colon\Sigma\rightarrow B$ be a Riemann surface over $B$ with a section $\sigma$. An irregular type on $B$ with pole at $\sigma$ of order at most $k$ is a global section of the sheaf $\mathfrak{t}\otimes_{\mathbf{C}}\mathscr{T}_{\Sigma,\sigma}^{\leq k}$.

### v2

@@ 2.2.3. Properties of $\mathscr{T}_{\Sigma,\sigma}^{\leq k}$
For $k=0$ the sheaf $\mathscr{T}_{\Sigma,\sigma}^{\leq k}$ is the zero sheaf; let us now assume that $k\geq 1$. For every $x\in\Sigma\smallsetminus\sigma(B)$ there is an open neighbourhood $U$ of $x$ which does not intersect $\sigma(B)$, hence $\mathcal{I}_{\sigma}|U=\mathcal{O}_{\Sigma}|U$. It follows that $\mathscr{T}_{\Sigma,\sigma}^{\leq k}$ is a sheaf supported on $\sigma(B)$ and annihilated by $\mathcal{I}_{\sigma}^{k}$, so it is the pushforward of a unique sheaf, denoted by $\mathscr{T}_{\sigma}^{\leq k}$, on the complex analytic space $(\sigma(B),\mathcal{O}_{\Sigma}/\mathcal{I}_{\sigma}^{k})$. In particular, even though $\mathscr{T}_{\Sigma,\sigma}^{\leq k}$ was defined as a quotient of a sheaf of meromorphic functions on the whole $\Sigma$, it only carries information about germs of such functions around $\sigma(B)$. Precisely, if $V\subset U\subset\Sigma$ are two opens having the same intersection with $\sigma(B)$, then $\mathscr{T}_{\Sigma,\sigma}^{\leq k}(U)=\mathscr{T}_{\Sigma,\sigma}^{\leq k}(V)$.
We will denote by $B^{(k)}$ the analytic space $(\sigma(B),\mathcal{O}_{\Sigma}/\mathcal{I}_{\sigma}^{k})$, and by $\pi^{(k)}\colon B^{(k)}\rightarrow B$ the composition of $\pi$ and the inclusion of $B^{(k)}$ in $\Sigma$.
@@
The sheaf $\pi_{*}\mathscr{T}_{\Sigma,\sigma}^{\leq k}$ coincides with the pushforward of $\mathscr{T}_{\sigma}^{\leq k}$ via the map $\pi^{(k)}\colon B^{(k)}\rightarrow B$. Choosing local coordinates around a point of $\sigma(B)$ as in section 2.1.2, we see that we can write locally $\mathscr{T}_{\Sigma,\sigma}^{\leq k}=z_{n+1}^{-k}\mathcal{O}_{\Sigma}/\mathcal{O}_{\Sigma}$, hence $\mathscr{T}_{\sigma}^{\leq k}$ is locally free of rank one.
On the other hand, we claim that $\pi^{(k)}_{*}\mathcal{O}_{B^{(k)}}$ is a locally free $\mathcal{O}_{B}$-module of rank $k$. To prove this, we may, and will, replace $B$ by an open neighbourhood of any of its points, and work with the same local coordinates as before. The section $\sigma\colon B\rightarrow\Sigma$ factors through $B^{(k)}$, inducing the natural projection $\mathcal{O}_{\Sigma}/z_{n+1}^{k}\mathcal{O}_{\Sigma}\rightarrow\mathcal{O}_{\Sigma}/z_{n+1}\mathcal{O}_{\Sigma}\simeq\mathcal{O}_{B}$. The map $\pi^{(k)}$ induces on global sections a section of this projection; hence, the claim follows from Lemma 2.2.5 just below. ∎
@@
Let $A$ be a commutative ring with unit, and $0\neq f\in A$ a nonzerodivisor which is not a unit. Let $k\geq 1$ be an integer and $s\colon A/(f)\rightarrow A/(f^{k})$ a ring morphism which is a section of the projection $q\colon A/(f^{k})\rightarrow A/(f)$. The map $s$ endows $A/(f^{k})$ with the structure of a free $A/(f)$-module of rank $k$ with basis $1,f,\ldots,f^{k-1}$.
@@
By induction on $k$; for $k=1$ the claim is clear, so let us suppose that $k\geq 2$.
Take $a\in A$ and let $\alpha_{0}=q(a)$. Then $q(a-s(\alpha_{0}))=q(a)-q\circ s\circ q(a)=0$, so we can write $a=s(\alpha_{0})+fa_{1}$ for some $a_{1}\in A/(f^{k})$. Assume that $\alpha_{0}^{\prime}\in A/(f),a_{1}^{\prime}\in A/(f^{k})$ also satisfy $a=s(\alpha_{0}^{\prime})+fa_{1}^{\prime}$. Then $s(\alpha_{0}-\alpha_{0}^{\prime})=f(a_{1}^{\prime}-a_{1})$; taking the image via $q$ we find that $\alpha_{0}=\alpha_{0}^{\prime}$, hence $f(a_{1}^{\prime}-a_{1})=0\in A/(f^{k})$. As $f$ is not a zero divisor, we deduce that $a_{1}^{\prime}-a_{1}\in(f^{k-1})$. Therefore, we have proved that $\alpha_{0}$ is unique, and $a_{1}$ is unique in $A/(f^{k-1})$, which implies the statement by induction. ∎
@@ 2.2.6. Functoriality
For $k\geq 0$, let us set $\mathscr{T}_{|B}^{\leq k}=\pi_{*}\mathscr{T}_{\Sigma,\sigma}^{\leq k}$. Let $\varphi\colon B^{\prime}\rightarrow B$ be a holomorphic map, $\pi^{\prime}\colon\Sigma^{\prime}\rightarrow B^{\prime}$ the base change of $\pi$, and $\sigma^{\prime}\colon B^{\prime}\rightarrow\Sigma^{\prime}$ the pullback of $\sigma$. We have a commutative diagram
@@
and the square with arrows $\sigma^{\prime},\sigma,\varphi$ and $\psi$ is Cartesian, hence $\sigma^{\prime}(B^{\prime})\subset\Sigma^{\prime}$ is cut out by the ideal sheaf image of $\psi^{*}\mathcal{I}_{\sigma}$. In other words, we have $\sigma^{\prime}_{*}\mathcal{O}_{B^{\prime}}=\mathcal{O}_{\Sigma^{\prime}}/\psi^{*}\mathcal{I}_{\sigma}=\psi^{*}\sigma_{*}\mathcal{O}_{B}$.
@@
The following identities hold:
@@
    $\psi^{*}\mathcal{I}_{\sigma}=\mathcal{I}_{\sigma^{\prime}}$;
@@
    $\psi^{*}\mathcal{O}_{\Sigma}(\sigma)=\mathcal{O}_{\Sigma^{\prime}}(\sigma^{\prime})$;
@@
    Pulling back the exact sequence $0\rightarrow\mathcal{I}_{\sigma}\rightarrow\mathcal{O}_{\Sigma}\rightarrow\sigma_{*}\mathcal{O}_{B}\rightarrow 0$ via $\psi$, and using that $\psi^{*}\sigma_{*}\mathcal{O}_{B}=\sigma^{\prime}_{*}\mathcal{O}_{B^{\prime}}$, we get an exact sequence
@@
    Pulling back the identity $\mathcal{I}_{\sigma}\otimes_{\mathcal{O}_{\Sigma}}\mathcal{O}_{\Sigma}(\sigma)\simeq\mathcal{O}_{\Sigma}$ and using (1) we deduce that $\psi^{*}(\mathcal{O}_{\Sigma}(\sigma))$ is the inverse of $\mathcal{I}_{\sigma^{\prime}}$, hence it coincides with $\mathcal{O}_{\Sigma^{\prime}}(\sigma^{\prime})$.
@@
    The $k$-fold tensor product of the isomorphism in (2) yields an isomorphism $\psi^{*}\mathcal{O}_{\Sigma}(k\sigma)=\mathcal{O}_{\Sigma^{\prime}}(k\sigma^{\prime})$, which implies the claim by right-exactness of pullback. ∎
By adjunction we have a natural transformation $\mathrm{Id}\rightarrow\psi_{*}\psi^{*}$, inducing $\pi_{*}\rightarrow\pi_{*}\psi_{*}\psi^{*}$, so $\pi_{*}\rightarrow\varphi_{*}\pi^{\prime}_{*}\psi^{*}$ and finally $\varphi^{*}\pi_{*}\rightarrow\pi^{\prime}_{*}\psi^{*}$. Applying these functors to $\mathscr{T}_{\Sigma,\sigma}^{\leq k}$ and using Lemma 2.2.7 we get a map
(2.2.7.1) $\varphi^{*}\mathscr{T}_{|B}^{\leq k}\rightarrow\mathscr{T}_{|B^{\prime}}^{\leq k}.$
@@
The map (2.2.7.1) is an isomorphism.
@@
By [Sta18, Tag 01HQ] and Lemma 2.2.7 there is a unique map $\gamma\colon B^{\prime(k)}\rightarrow B^{(k)}$ making the top square of the diagram in section 2.2.6 Cartesian.
Then we claim that the following diagram is Cartesian as well:
@@ 2.3. Irregular types
Let $\pi\colon\Sigma\rightarrow B$ be a Riemann surface over $B$ with a section $\sigma$, and $k\geq 0$ an integer. We define an irregular type (on $B$) with pole at $\sigma$ of order at most $k$ to be a global section of the sheaf $\mathfrak{t}\otimes_{\mathbf{C}}\mathscr{T}_{\Sigma,\sigma}^{\leq k}$.
2.3.1. Irregular type of a connection on a framed bundle
Assume in this subsection that $B$ is a point, and denote the image of $\sigma$ by $a$. Let $K$ be the fraction field of the completion of the local ring $\mathcal{O}_{\Sigma,a}$, with its valuation $v:K\rightarrow\mathbf{Z}\cup\{+\infty\}$. For $l\in\mathbf{Z}$, let $K_{\geq l}$ be the set of elements $x\in K$ such that $v(x)\geq l$. We will denote the valuation ring $K_{\geq 0}$ by $\mathcal{O}_{K}$, and the maximal ideal $K_{\geq 1}$ of $\mathcal{O}_{K}$ by $\mathfrak{m}$. With this notation, an irregular type with pole at $\sigma$ of order at most $k$ is an element of $\mathfrak{t}\otimes_{\mathbf{C}}(K_{\geq-k}/\mathcal{O}_{K})$. Let us recall how, for $\mathfrak{g}=\mathfrak{gl}_{r}$ and $\mathfrak{t}$ the Cartan algebra of diagonal matrices, one can attach to certain “framed” vector bundles with meromorphic connection on $\Sigma$ a unique irregular type. This construction is not used elsewhere in this document, but we include it because it is the main motivation for the definition of irregular type. See [Boa02, §2] for a similar discussion in the generic case, for general $\mathfrak{g}$.
Let $\mathcal{V}$ be a vector bundle of rank $r$ on $\Sigma$, and $\nabla$ a meromorphic connection on $\mathcal{V}$ with a pole at $a$ of order at most $k+1$. Looking at the stalk at $a$, and further pulling back to the completed local ring, we obtain a free $\mathcal{O}_{K}$-module $V$ of rank $r$ together with a meromorphic connection, abusively still denoted by the same symbol,
$\nabla:V\rightarrow\hat{\Omega}^{1}_{K}\otimes_{\mathcal{O}_{K}}V,$
where $\hat{\Omega}^{1}_{K}$ is the module of continuous Kähler differentials of $K/\mathbf{C}$. Choosing a basis $(e_{1},\ldots,e_{r})$ of $V$ we obtain an isomorphism of $\mathcal{O}_{K}$-modules $V\simeq\mathcal{O}_{K}^{r}$, hence there exists a unique $r\times r$ matrix $M$ with coefficients $\omega_{ij}\in\hat{\Omega}^{1}_{K}$ such that $\nabla(e_{i})=\sum_{j=1}^{r}\omega_{ij}\otimes e_{j}$ for $1\leq i\leq r$. Changing basis - replacing the column matrix $E$ with entries $e_{1},\ldots,e_{r}$ by $gE$ for $g\in\mathrm{GL}_{r}(\mathcal{O}_{K})$ - results in replacing $M$ by $gMg^{-1}+\mathrm{d}gg^{-1}$.
Let $\hat{\Omega}^{1}_{\mathcal{O}_{K}}$ be the module of continuous Kähler differentials of $\mathcal{O}_{K}/\mathbf{C}$, $\hat{\Omega}^{1}_{K,\geq-1}=\hat{\Omega}^{1}_{\mathcal{O}_{K}}\otimes_{\mathcal{O}_{K}}K_{\geq-1}$ and $\hat{\Omega}^{1}_{K,<-1}=\hat{\Omega}^{1}_{K}/\hat{\Omega}^{1}_{K,\geq-1}$. For $\omega\in\hat{\Omega}^{1}_{K}$, we denote by $\bar{\omega}$ its image in $\hat{\Omega}^{1}_{K,<-1}$; let $\bar{M}$ be the matrix with entries $\bar{\omega}_{ij}$. We say that $\nabla$ is untwisted if there exists a basis of $V$ such that $\bar{M}$ is diagonal (for a general meromorphic connection, one needs to replace $K$ by a finite extension in order to obtain a diagonal $\bar{M}$, see [Tur55, Theorem 1]). We will explain how to attach to such a $\nabla$ an element of $\mathfrak{t}\otimes(K_{\geq-k}/\mathcal{O}_{K})$ which only depends on the reduction of $(e_{1},\ldots,e_{r})$ modulo $\mathfrak{m}$, i.e., on the choice of a suitable basis of the fibre of $\mathcal{V}$ at $a$.
First of all, note that the derivation $d\colon K\rightarrow\hat{\Omega}^{1}_{K}$ induces an isomorphism $K/K_{\geq 0}\xrightarrow{\sim}\hat{\Omega}^{1}_{K,<-1}$. Therefore, given an untwisted connection $\nabla$ with a pole at $a$ of order at most $k+1$ and a basis $(e_{1},\ldots,e_{r})$ of $V$ such that $\bar{M}$ is diagonal, we obtain a unique element $IT(\nabla,(e_{1},\ldots,e_{r}))\in\mathfrak{t}\otimes(K_{\geq-k}/\mathcal{O}_{K})$ mapping to $\bar{M}\in\mathfrak{t}\otimes_{\mathbf{C}}\hat{\Omega}^{1}_{K,<-1}$. It remains to show that $IT(\nabla,(e_{1},\ldots,e_{r}))$ only depends on the reduction of $(e_{1},\ldots,e_{r})$ modulo $\mathfrak{m}$. In other words, if $g\in\mathrm{GL}_{r}(\mathcal{O}_{K})$ projects to the identity in $\mathrm{GL}_{r}(\mathbf{C})$ and the image $\bar{M}^{\prime}$ of $M^{\prime}=gMg^{-1}$ in $\mathfrak{g}\otimes_{\mathbf{C}}\hat{\Omega}^{1}_{K,<-1}$ is diagonal, we need to prove that $\bar{M}^{\prime}=\bar{M}$.
To show this, let us fix an isomorphism $\mathcal{O}_{K}\simeq\mathbf{C}[[z]]$, yielding an isomorphisms $\hat{\Omega}^{1}_{K}\simeq\mathbf{C}((z))\mathrm{d}z$. This allows us to see $M$ and $M^{\prime}$ as matrices with coefficients in $\frac{1}{z^{k+1}}\mathbf{C}[[z]]$, and $g$ as an element of $GL_{r}(\mathbf{C}[[z]])$ congruent to the identity modulo $(z)$. Multiplying $M$ and $M^{\prime}$ by $z^{k+1}\mathrm{Id}$ turns them into matrices $N$ and $N^{\prime}$ with $\mathbf{C}[[z]]$-coefficients. Let $\bar{N}$ and $\bar{N}^{\prime}$ be the images of $N$ and $N^{\prime}$ in $A=\mathbf{C}[[z]]/(z^{k})$; we need to show that $\bar{N}=\bar{N}^{\prime}$. This follows for instance from [CFRW, Lemma 8.1.2], which proves a more general statement, valid for arbitrary $\mathfrak{g}$; let us give a simple argument in the case of $\mathfrak{gl}_{r}$. Let $(v_{1},\ldots,v_{r})$ be the canonical basis of $A^{r}$. By assumption, there exist $\lambda_{1},\ldots,\lambda_{r}\in A$ such that $\bar{N}v_{i}=\lambda_{i}v_{i}$ for $1\leq i\leq r$, hence $\bar{N}^{\prime}(gv_{i})=\lambda_{i}\cdot gv_{i}$. On the other hand, we are assuming that $\bar{N}^{\prime}v_{i}=\gamma_{i}v_{i}$ for some $\gamma_{1},\ldots,\gamma_{r}\in A$. Writing
$gv_{i}=\sum_{j=1}^{r}a_{i,j}v_{j}\text{ with }a_{i,j}\in A$
we obtain $\bar{N}^{\prime}(gv_{i})=\sum_{j=1}^{r}a_{i,j}\gamma_{j}v_{j}=\sum_{j=1}^{r}a_{i,j}\lambda_{i}v_{j}$. Finally, the vectors $g(v_{i})$ and $v_{i}$ have the same image in $(A/(z))^{r}$, hence $a_{i,i}$ belongs to $A^{\times}$ and $\gamma_{i}=\lambda_{i}$.